OpenStudy (anonymous):

If the area of a circle is approximately 113 square yards, about how long is the radius? (Use 3.14 for π.)

OpenStudy (anonymous):

HELLP

OpenStudy (trojanpoem):

Bro, all you have to know is the relation of circle area. area of circle = pi (r)^2 area = 113 ( given) pi = 3.14 (given) 113 = 3.14 r^2 , r = 5.999 Advice : Next time it will be easier to use google. All you have to do is to write "Circle area relation".